OpenStudy (anonymous):

how do i identify the horizontal translation of the parent function, f(x)= x squared y=(x-4) squared

OpenStudy (anonymous):

\[f(x)=a(x-h)^{2}+k\] the horizontal shift is the value of x that makes (x-h) = 0

OpenStudy (anonymous):

for yours, you have \[f(x)=(x-4)^{2}\] what value of x makes (x-4) = 0?

OpenStudy (anonymous):

so basically plug in f(x) for y?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

if \[f(x) = x^{2}\] is the parent function and \[g(x)=(x-4)^{2}\] is the transformed function then g(x) is a horizontal shift of f(x). The amount of horzontal shift is given by finding what value of x makes the (x-4) term equal 0.